# Eigenvalue degeneracy in sparse random matrices

### v1

@@ Abstract
eigenvalue degeneracy, random matrices, sparse matrices, random graphs
MSC Classification: 60B20, 05C80
@@ 1 Introduction
This theorem appeared in the paper[2] by von Neumann and Wigner published in 1929. Let us explain it for a hermitian matrix. Suppose that an $N\times N$ hermitian matrix has $f$ distinct eigenvalues and the multiplicity of each distinct eigenvalue is $g_{1},\cdots,g_{f}$. Then such a hermitian matrix can be written in terms of $N^{2}+f-g_{1}^{2}-g_{2}^{2}-\cdots-g_{f}^{2}$ parameters. In a special case $g_{1}=g_{2}=\cdots=g_{f-1}=1$ with $g_{f}=2$, we need $N^{2}-3$ parameters. As a hermitian matrix in general is written in terms of $N^{2}$ parameters, the codimension is $3$. Therefore, if the probability measure of the whole space of hermitian matrices is absolutely continuous with respect to the $N^{2}$ dimensional Lebesgue measure, the degeneracy probability of the matrix eigenvalues is zero.
@@
In §3, matrix models related to bipartite graphs are introduced. In these models, if the $j$-th and $\ell$-th vertices are not connected by an edge in the bipartite graph, then the $(j,\ell)$ entry of the corresponding matrix is fixed to zero. The concept of perfect matching is borrowed from the theory of bipartite graphs and used in the study of degeneracy. If a bipartite graph has a perfect matching, the eigenvalues of at least one of the corresponding matrices do not degenerate. Such an analysis using the concept of graph theory is also applied to symmetric matrices.
In §4, employing the results of §2 and §3 together, we study the eigenvalue degeneracy probability for $N\times N$ sparse random matrices. The matrix entries are supposed to be continuously distributed expect at the origin. In order to evaluate the degeneracy probability, we refer to a theory by Erdös and Rényi concerning the perfect matching probability of random bipartite graphs. As a result, the degeneracy probability is asymptotically evaluated in the limit $N\rightarrow\infty$, and a positive degeneracy probability is worked out. Accumulation of the eigenvalues to the origin, due to the discontinuity of the matrix entry distribution, causes the degeneracy. In §5, we consider the degeneracy probability for symmetric sparse random matrices, and again find a positive asymptotic limit.
@@ 2 Eigenvalue degeneracy probability for continuous distributions
Matrix eigenvalues are roots of the characteristic equation. In order to study the root degeneracy, we first introduce the concept of a discriminant. Using discriminants, we show a sufficient condition for the eigenvalue degeneracy probability of random matrices to be zero or one.
@@
Because of Lemma 2.2, a polynomial $p(\lambda)$ has a multiple root, if and only if the discriminant $\Delta[p(\lambda)]$ is zero.
@@ 3.1 Asymmetric matrices
Theorem 3.6 (Hall’s marriage theorem)
@@ 4 Eigenvalue degeneracy probability for random matrices
Thus, as a simpler problem, let us first consider the probability for a random graph to have a perfect matching. For that problem, there is a known result explained below as Theorem 4.2.
@@ 4.1 The probability for a random graph to have a perfect matching
In this subsection, we introduce the following theorem for the asymptotic probability for a random bipartite graph to have a perfect matching. This theorem was originally proved by Erdös and Rényi[9]. It is useful in the study of asymptotic degeneracy probability for the eigenvalues of the corresponding random matrices,
Theorem 4.2
@@
In the remaining part of this subsection, we give a step-by-step proof of Theorem 4.2 in reference to Frieze and Karoński’s book[10]. Intermediate results are also useful in the study of eigenvalue degeneracy. Let us begin with a definition and a lemma.
@@
and Lemma 4.4, we further find
@@
holds due to Lemma 4.5. Taking the limit $n\rightarrow\infty$, we find
@@
Lemma 4.9
Let us come back to the notations of Theorem 4.2 and again consider a random bipartite graph $G_{n}=({\cal V}_{1},{\cal V}_{2},{\cal E})$ with ${\cal E}\subset{\cal V}_{1}\times{\cal V}_{2}$ and ${\cal V}_{1}=\#{\cal V}_{2}=n$. Each $\langle j,\ell\rangle\in{\cal V}_{1}\times{\cal V}_{2}$ ($j,\ell=1,2,\cdots,n$) independently satisfies $\langle j,\ell\rangle\in{\cal E}$ with probability $p$. The probability parameter $p$ depends on $n$ as
@@
It follows from Lemma 4.7 that
@@
Now we can utilize Corollary 4.6 to find
@@
In order to prove Theorem 4.2, we need to show that the probability for a random bipartite graph $G$ to have neither isolated points nor perfect matchings converges to zero.
When a bipartite graph $G=({\cal V}_{1},{\cal V}_{2},{\cal E})$ has no perfect matchings, Theorem 3.6 (Hall’s marriage theorem) means that
@@
The same theorem can also be applied to the case when ${\cal V}_{1}$ and ${\cal V}_{2}$ are interchanged and consequently
@@
As for such sets $A_{1}$ and $A_{2}$, we have the following lemma.
Lemma 4.10
@@
Due to Lemma 4.10, if a bipartite graph $G$ has no perfect matchings, then it has at least one isolated point or satisfies the following condition.
@@
holds. We can see from Lemma 4.9, Lemma 4.12 and the squeeze theorem that
@@ 4.2 Eigenvalue degeneracy probability for sparse random matrices
Because of Lemma 4.9, the first term of the RHS satisfies
@@
We can see from Lemma 4.9 that the corresponding asymptotic probability is
@@
In the proof of Lemma 4.9, we showed that $\mathbb{E}[X_{N}]<\infty$. Moreover, as
@@
We can again see from the proof of Lemma 4.9 that $\mathbb{E}[X_{N}]<\infty$. Since p(N)=p(N1)+o(1/N)p(N)=p(N-1)+o(1/N), Lemma 4.12 leads to
@@
If $X_{N}=0$, in order to make Condition 4.1 unsatisfied, Condition 4.11 must be satisfied. It follows from Lemma 4.12 that
@@ 5 Eigenvalue degeneracy in symmetric sparse random matrices
It follows from Lemma 4.6 that
@@
which also holds for $k=0$. Thus one can see from Corollary 4.5 that
@@
We can prove (1), (2) and (3) in the same way as Lemma 4.10. A proof of (4) is given below.

### v2

@@ Abstract
eigenvalue degeneracy; random matrices; sparse matrices; random graphs.
Mathematics Subject Classification: 60B20, 05C80
@@ 1 Introduction
This proposition appeared in the paper[2] by von Neumann and Wigner published in 1929. Let us explain it for a hermitian matrix. Suppose that an $N\times N$ hermitian matrix has $f$ distinct eigenvalues and the multiplicity of each distinct eigenvalue is $g_{1},\cdots,g_{f}$. Then such a hermitian matrix can be written in terms of $N^{2}+f-g_{1}^{2}-g_{2}^{2}-\cdots-g_{f}^{2}$ parameters. In a special case $g_{1}=g_{2}=\cdots=g_{f-1}=1$ with $g_{f}=2$, we need $N^{2}-3$ parameters. As a hermitian matrix in general is written in terms of $N^{2}$ parameters, the codimension is $3$. Therefore, if the probability measure of the whole space of hermitian matrices is absolutely continuous with respect to the $N^{2}$ dimensional Lebesgue measure, the degeneracy probability of the matrix eigenvalues is zero.
@@
In §3, matrix models related to bipartite graphs are introduced. In these models, if the $j$-th and $\ell$-th vertices are not connected by an edge in the bipartite graph, then the $(j,\ell)$ entry of the corresponding matrix is fixed to zero. The concept of perfect matching is borrowed from the theory of bipartite graphs and used in the study of degeneracy. If a bipartite graph has a perfect matching, the eigenvalues of at least one of the corresponding matrices do not degenerate. Such an analysis using the concept of graph theory is also applied to symmetric matrices. Moreover, suppose that $\phi$ is an $N\times N$ matrix with elements
$\phi_{j\ell}=\begin{cases}x_{j\ell},&(j,\ell)\in\Lambda,\\ 0,&\text{otherwise}.\end{cases}$ (1.1)
Here $x_{j\ell}\in\mathbb{R}\text{ or }\mathbb{C}$ and $\Lambda\subset\{(j,\ell)\mid 1\leq j,\ell\leq N\}$. Then, if the cardinality of $\Lambda$ is larger than $N^{2}-2N+1$, a family of $x_{j\ell}$ ($(j,\ell)\in\Lambda$) exists, so that the eigenvalues of $\phi$ do not degenerate.
In §4, employing the results of §2 and §3 together, we study the eigenvalue degeneracy probability for $N\times N$ sparse random matrices. Let us suppose that random variables $Y_{j\ell}$ ($j,\ell=1,2,\cdots,N)$ independently obey Bernoulli distribution with a parameter $p$. That is, $Y_{j\ell}$ takes value $1$ with probability $p$ and 0 with probability 1p1-p. Here $p$ depends on $N$ as
with a real $c$ fixed. Moreover we assume that continuous random variables $Z_{j\ell}$ ($j,\ell=1,2,\cdots,N$) are independently distributed and consider the eigenvalue degeneracy probability of an $N\times N$ sparse random matrix $(Y_{j\ell}Z_{j\ell})_{j,\ell=1,2,\cdots,N}$. The random matrix entries are continuously distributed except at the origin. In order to asymptotically evaluate the degeneracy probability in the limit $N\rightarrow\infty$, we refer to a theory by Erdös and Rényi concerning the perfect matching probability of random bipartite graphs. As a result, a positive asymptotic limit
$1-e^{-\lambda}-\lambda e^{-\lambda}$ (1.3)
is worked out ($\lambda=2e^{-c}$) for the degeneracy probability. Accumulation of the eigenvalues to the origin, due to the discontinuity of the matrix entry distribution, causes the degeneracy. In §5, we consider the degeneracy probability for symmetric sparse random matrices, and again find a positive asymptotic limit.
@@ 2 Eigenvalue degeneracy probability for continuous distributions
Matrix eigenvalues are roots of the characteristic equation. In order to study the root degeneracy, we first introduce the concept of a discriminant. The discriminant detects multiple roots and is expressed as the resultant of a polynomial and its derivative, which is given by the determinant of the Sylvester matrix (see, e.g., Gelfand-Kapranov-Zelevinsky [6], Ch. 12, §1). Using discriminants, we show a sufficient condition for the eigenvalue degeneracy probability of random matrices to be zero or one.
@@
Because of Proposition 2.2, a polynomial $p(\lambda)$ has a multiple root, if and only if the discriminant $\Delta[p(\lambda)]$ is zero.
@@ 3.1 Asymmetric matrices
Proposition 3.6 (Hall’s marriage theorem)
@@ 4 Eigenvalue degeneracy probability for random matrices
Thus, as a simpler problem, let us first consider the probability for a random graph to have a perfect matching. For that problem, there is a known result explained below as Proposition 4.2.
@@ 4.1 The probability for a random graph to have a perfect matching
In this subsection, we introduce the following proposition for the asymptotic probability for a random bipartite graph to have a perfect matching. This proposition was originally proved by Erdös and Rényi[10]. It is useful in the study of asymptotic degeneracy probability for the eigenvalues of the corresponding random matrices,
Proposition 4.2
@@
In the remaining part of this subsection, we give a step-by-step proof of Proposition 4.2 in reference to Frieze and Karoński’s book[11]. Intermediate results are also useful in the study of eigenvalue degeneracy. Let us begin with a definition and a proposition.
@@
and Proposition 4.4, we further find
@@
holds due to Proposition 4.5. Taking the limit $n\rightarrow\infty$, we find
@@
Proposition 4.9
Let us come back to the notations of Proposition 4.2 and again consider a random bipartite graph $G_{n}=({\cal V}_{1},{\cal V}_{2},{\cal E})$ with ${\cal E}\subset{\cal V}_{1}\times{\cal V}_{2}$ and ${\cal V}_{1}=\#{\cal V}_{2}=n$. Each $\langle j,\ell\rangle\in{\cal V}_{1}\times{\cal V}_{2}$ ($j,\ell=1,2,\cdots,n$) independently satisfies $\langle j,\ell\rangle\in{\cal E}$ with probability $p$. The probability parameter $p$ depends on $n$ as
@@
It follows from Proposition 4.7 that
@@
Now we can utilize Proposition 4.6 to find
@@
In order to prove Proposition 4.2, we need to show that the probability for a random bipartite graph $G$ to have neither isolated points nor perfect matchings converges to zero.
When a bipartite graph $G=({\cal V}_{1},{\cal V}_{2},{\cal E})$ has no perfect matchings, Proposition 3.6 (Hall’s marriage theorem) means that
@@
The same proposition can also be applied to the case when ${\cal V}_{1}$ and ${\cal V}_{2}$ are interchanged and consequently
@@
As for such sets $A_{1}$ and $A_{2}$, we have the following proposition.
Proposition 4.10
@@
Due to Proposition 4.10, if a bipartite graph $G$ has no perfect matchings, then it has at least one isolated point or satisfies the following condition.
@@
holds. We can see from Proposition 4.9, Proposition 4.12 and the squeeze theorem that
@@ 4.2 Eigenvalue degeneracy probability for sparse random matrices
Because of Proposition 4.9, the first term of the RHS satisfies
@@
We can see from Proposition 4.9 that the corresponding asymptotic probability is
@@
In the proof of Proposition 4.9, we showed that $\mathbb{E}[X_{N}]<\infty$. Moreover, as
@@
We can again see from the proof of Proposition 4.9 that $\mathbb{E}[X_{N}]<\infty$. Since p(N)=p(N1)+o(1/N)p(N)=p(N-1)+o(1/N), Proposition 4.12 leads to
@@
If $X_{N}=0$, in order to make Condition 4.1 unsatisfied, Condition 4.11 must be satisfied. It follows from Proposition 4.12 that
@@ 5 Eigenvalue degeneracy in symmetric sparse random matrices
It follows from Proposition 4.7 that
@@
which also holds for $k=0$. Thus one can see from Proposition 4.6 that
@@
We can prove (1), (2) and (3) in the same way as Proposition 4.10. A proof of (4) is given below.